20 Liters of milk costs 40 dollars.
How much does a lt. of milk costs ?
2 dollars or 0,5 cents ?

Answer by stanbon(75887)

20 Liters of milk costs 40 dollars.
How much does a lt. of milk costs ?
2 dollars or 0,5 cents ?
--------
x/1 lt = $40/20 lt
------------
x = 40/20 = $2.00
--------------------
Cheers,
Stan H.
--------------------